# RoHOI: Robustness Benchmark for Human-Object Interaction Detection

### v2

@@ Abstract
Human-Object Interaction (HOI) detection is crucial for robot-human assistance, enabling context-aware support. However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate prediction. To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges. Despite advances, current models struggle with environmental variability, occlusions, and noise. Our benchmark, RoHOI, includes 20 corruption types based on the HICO-DET and V-COCO datasets and a new robustness-focused metric. We systematically analyze existing models in the HOI field, revealing significant performance drops under corruptions. To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, thus dynamically adjusting the model’s optimization to enhance robust feature learning. Extensive experiments show that our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection. Benchmarks, datasets, and code will be made publicly available at https://github.com/Kratos-Wen/RoHOI.
 [^1] The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation). This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026).
Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups. Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions. These corruptions simulate realistic semantic and structural disturbances specifically encountered in practical application.
Introduction
@@
Human-Object Interaction (HOI) detection aims to identify and interpret interactions between humans and objects (Kim et al. 2021; Hou et al. 2021a; Zhang et al. 2022; Tamura, Ohashi, and Yoshinaga 2021; Kim et al. 2022; Hou et al. 2021b). It serves many critical applications, including autonomous driving (Martin et al. 2019; Peng et al. 2022), robotics (Xu et al. 2020), video surveillance (Dogariu et al. 2020; Prest, Ferrari, and Schmid 2013), and augmented reality (Zhang, Cui, and Fang 2025), where a fine-grained understanding of human behavior and object handling is paramount (Gupta and Malik 2015).
@@
In recent years, transformer-based models (Zhang et al. 2022; Tu et al. 2023; Zou et al. 2021) and Vision-Language Models (VLMs) (Cao et al. 2023; Gao et al. 2024; Chen et al. 2024a), have achieved impressive success in the HOI detection field. These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information.
@@
Despite these advancements, existing models remain sensitive to real-life corruptions, particularly semantic distortions like occlusions and perspective shifts, as well as environmental variations. Even minor disturbances may pose significant risks in critical applications like autonomous driving and robotics. Enhancing robustness under these conditions remains a crucial but unresolved challenge. Existing HOI detection methods (Zhang et al. 2022; Tu et al. 2023; Cao et al. 2023) had their performance evaluated predominantly under ideal conditions. Thus, a dedicated robustness benchmark tailored explicitly to HOI detection, covering both pixel-level and semantic-level corruptions, is urgently required to put their resilience to the test.
@@
In this paper, we introduce the Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse forms of data corruption. These were designed to simulate real-world conditions. The HICO-DET (Chao et al. 2018) and V-COCO (Gupta and Malik 2015) datasets served as starting point. We include robustness-focused metrics, such as the Mean Robustness Index (MRI) (Hendrycks and Dietterich 2019) and our novel Composite Robustness Index (CRI). Through our evaluations of existing HOI detection methods on the newly proposed robustness test set, we observe that their performance is less than optimal. This finding underscores the need to develop more robust HOI detection models.
@@
To mitigate the issue of performance degradation, we propose our Semantic-Aware Masking-based Progressive Learning (SAMPL) method. Specifically, we utilize SAM (Kirillov et al. 2023) to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying size ratios. This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process.
@@
Related Work
@@
Human Object Interaction Detection. Human-object interaction (HOI) detection (Yao and Fei-Fei 2010; Fang et al. 2018; Gupta, Schwing, and Hoiem 2019; Li et al. 2019; Wan et al. 2019; Park et al. 2018; Xu et al. 2020; Yu, Liu, and Odobez 2018; Zhang et al. 2015; Zhou et al. 2020) aims to automatically identify and interpret relationships between humans and objects, typically divided into two categories: two-stage and one-stage approaches.
@@
Two-stage HOI detection methods (Qi et al. 2018; Xu et al. 2019; Liang et al. 2021; Zhou and Chi 2019) first detect humans and objects independently before classifying interactions. Early CNN-based approaches (Xu et al. 2019; Zhou and Chi 2019; Qi et al. 2018) mainly utilized spatial contexts for interaction modeling. Later graph-based methods (Zhou and Chi 2019; Xu et al. 2019; Ulutan, Iftekhar, and Manjunath 2020; Wang et al. 2021; Zhang, Campbell, and Gould 2021) improved relational reasoning explicitly. Subsequent methods further enhanced precision via pose-aware representations (Wan et al. 2019), transformer-based unpaired modeling (Zhang, Campbell, and Gould 2022), and integration-decomposition strategies (Li et al. 2020).
One-stage HOI detection methods (Yuan et al. 2023; Zhang et al. 2023) achieve efficiency by directly predicting human-object interactions end-to-end. Early methods exploited parallel point detection (Liao et al. 2020) and union-level modeling (Kim et al. 2020), while subsequent approaches refined interaction localization via adaptive glance-and-gaze mechanisms (Zhong et al. 2021). Transformer-based methods enhanced relational context modeling (Kim, Jung, and Cho 2023) and language-image fusion (RLIPv2)(Yuan et al. 2023), significantly improving performance. Recently, vision-language models (VLMs) were leveraged to boost generalization across unseen interactions (VLM-HOI(Kang et al. 2024)) and zero-shot triplet predictions (EZ-HOI (Lei, Wang, and Tan 2024)). Robustness-specific methods further integrated uncertainty estimation (UAHOI (Chen, Chen, and Yang 2024)) to improve stability under perturbations. Nevertheless, all these approaches primarily evaluate performance on clean data, neglecting real-world corruptions. To address this gap, we introduce RoHOI, the first HOI robustness benchmark encompassing $20$ test-time corruptions.
@@
Visual Robustness Benchmarks. Robustness benchmarks have been widely established to assess visual model resilience across various tasks (Chattopadhyay et al. 2021; Wang et al. 2024; Sarkar et al. 2024; Schiappa et al. 2024; An et al. 2024; Liu et al. 2024b; Yi et al. 2024; Kamann and Rother 2020; Ma et al. 2024b; Hao et al. 2024; Xie et al. 2025; Hao et al. 2025; Liu et al. 2024c; Michaelis et al. 2019; Zeng et al. 2024), including image classification (ARES-Bench (Liu et al. 2024a)), visual question answering (Ishmam et al. 2024), common corruption evaluation (ImageNet-C (Hendrycks and Dietterich 2019)), attribute editing robustness (ImageNet-E (Li et al. 2023)), one-shot skeleton-based action recognition under occlusion (Peng et al. 2023), and robust document layout analysis (RoDLA (Chen et al. 2024b)). However, robustness benchmarks explicitly designed for HOI detection have not been explored. In this work, we address the uncharted robust HOI detection area by constructing the new RoHOI benchmark and by putting forward the SAMPL method to achieve more reliable HOI detection performance.
Benchmark
@@
Corruptions. To rigorously assess the robustness of HOI detection, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under real-world corruptions. Existing benchmarks such as ImageNet-C (Hendrycks and Dietterich 2019) primarily cover low-level pixel disturbances like brightness or blur. In contrast, RoHOI explicitly includes semantic-level corruptions specifically tailored to challenge spatial and relational reasoning critical for HOI detection. RoHOI explicitly introduces 20 corruptions derived from HICO-DET and V-COCO images, categorized into four groups: optical system artifacts such as motion blur and glass blur; sensor, compression, and transmission distortions such as packet loss and JPEG artifacts; environmental effects including exposure imbalance and occlusion; and geometric or structural disruptions such as screen cracks, elastic deformation, and perspective distortion. The corruptions range from pixel-level noise to semantic-level challenges that affect spatial alignment and interaction reasoning. Each corruption is implemented with five severity levels to support detailed analysis under progressively more difficult conditions.
@@
Datasets. To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO. HICO-DET (Wang et al. 2020) contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories. Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios. V-COCO (Gupta and Malik 2015) is a subset of MS-COCO (Lin et al. 2014) with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes. Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions. Evaluating these datasets under corruption provides a systematic robustness assessment beyond clean-image performance.
@@
Baselines. To systematically evaluate the robustness of HOI detection models, we benchmark both two-stage and one-stage architectures, including transformer-based and vision-language pretraining approaches. Two-stage methods first detect objects before predicting interactions. We evaluate QPIC (Tamura, Ohashi, and Yoshinaga 2021) and UPT (Zhang, Campbell, and Gould 2022), both of which use object detectors to extract human-object proposals before interaction classification. We include CDN (Zhang et al. 2021), which refines interaction reasoning via a cascade structure, and DiffHOI (Yang et al. 2023), which leverages text-to-image diffusion priors to improve rare interaction recognition. One-stage methods directly predict interactions without explicit object detection, improving efficiency. We evaluate FGAHOI (Ma et al. 2024a), which enhances spatial alignment using fine-grained anchors, and QAHOI (Chen and Yanai 2023), which integrates query-based attention for interaction localization. Additionally, we include MUREN (Kim, Jung, and Cho 2023), which refines relational context modeling, and SOV-STG (Chen, Wang, and Yanai 2023), which optimizes decoder attention for verb-object association. GEN-VLKT (Liao et al. 2022) incorporates visual-language pretraining to improve interaction recognition. RLIPv2 (Yuan et al. 2023) incorporates region-language alignment for fine-grained interaction reasoning. Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning. Our benchmark systematically evaluates these models across diverse corruption types and severity levels, providing a comprehensive assessment of HOI detection robustness.
@@
Evaluation Metrics. We evaluate the clean test set of HICO-DET (Wang et al. 2020) using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare (<<10 instances), and non-rare ($\geq$10 instances) categories (Yuan et al. 2023). For the clean test set of V-COCO (Gupta and Malik 2015), we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$. To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI). MRI, following the definition in prior work (Hendrycks and Dietterich 2019), measures average performance under corruptions but overlooks stability across severity levels. To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
@@
$\text{MRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{1}{L_{c}}\sum_{l=1}^{L_{c}}mAP_{c,l}\right),$ (1)
where $C$ denotes the number of corruption types, $L_{c}$ represents the number of severity levels for the $c$-th corruption, and $mAP_{c,l}$ is the model’s performance at level $l$ of corruption $c$. While MRI provides an overall robustness estimate, it does not account for performance fluctuations across corruption levels.
Composite Robustness Index (CRI) addresses MRI’s limitations. It penalizes performance instability across corruption levels while normalizing against clean-set performance:
$\text{CRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{\overline{mAP}_{c}}{mAP_{clean}}\cdot\frac{1}{\log(1+\sigma_{c})+1}\right),$ (2)
where $\overline{mAP}_{c}$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels. The clean-set performance $mAP_{clean}$ is defined as mAP (Full) for HICO-DET and $\mathrm{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5. The penalty term $\frac{1}{\log(1+\sigma_{c})+1}$ reduces scores for models with high variance, ensuring that robustness accounts for both overall performance and stability.
By integrating absolute accuracy with degradation consistency, CRI favors models that maintain stable performance across corruption levels, mitigating the impact of abrupt performance drops. For completeness, we also report standard HICO-DET mAP and V-COCO role-based AP scores, following prior benchmarks (detailed in Supplementary).
Method
In this section, we present our novel Semantic-Aware Masking-based Progressive Learning (SAMPL), aiming to enhance model robustness by progressively guiding models from holistic feature reliance towards robust contextual inference. Unlike standard augmentation methods, SAMPL employs adaptive semantic masking guided by SAM (Kirillov et al. 2023), dynamically adjusting masking severity based on validation performance. This structured semantic-level masking strategy explicitly promotes robust contextual reasoning by training models to infer interactions from progressively restricted semantic regions, thus effectively improving general robustness against a variety of realistic disturbances. Moreover, SAMPL’s structured progressive curriculum inherently regularizes feature representation, effectively reducing overfitting to dataset biases, thereby further improving generalization under unforeseen corruptions.

### v3

@@ Abstract
Human-Object Interaction (HOI) detection is crucial for robot-human assistance, enabling context-aware support. However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate predictions. To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges. Despite advances, current models struggle with environmental variability, occlusions, and noise. Our benchmark, RoHOI, includes 20 corruption types based on the HICO-DET and V-COCO datasets and a new robustness-focused metric. We systematically analyze existing models in the HOI field, revealing significant performance drops under corruptions. To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, thus dynamically adjusting the model’s optimization to enhance robust feature learning. Extensive experiments show that our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection. Benchmarks, datasets, and code are available at https://github.com/KratosWen/RoHOI.
1 Introduction
@@
Human-Object Interaction (HOI) detection aims to identify and interpret interactions between humans and objects Kim et al. (2021); Hou et al. (2021a); Zhang et al. (2022b); Tamura et al. (2021); Kim et al. (2022); Hou et al. (2021b). It serves many critical applications, including autonomous driving Martin et al. (2019); Peng et al. (2022), robotics Xu et al. (2020), video surveillance Dogariu et al. (2020); Prest et al. (2013), and augmented reality Zhang et al. (2025), where a fine-grained understanding of human behavior and object handling is paramount Gupta & Malik (2015).
@@
In recent years, transformer-based models Zhang et al. (2022b); Tu et al. (2023); Zou et al. (2021) and Vision-Language Models (VLMs) Cao et al. (2023); Gao et al. (2024); Chen et al. (2024b), have achieved impressive success in the HOI detection field. These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information.
@@
Despite these advancements, existing models remain sensitive to real-life corruptions, particularly semantic distortions like occlusions and perspective shifts, as well as environmental variations. Even minor disturbances may pose significant risks in critical applications like autonomous driving and robotics. Enhancing robustness under these conditions remains a crucial but unresolved challenge. Existing HOI detection methods Zhang et al. (2022b); Tu et al. (2023); Cao et al. (2023) had their performance evaluated predominantly under ideal conditions. Thus, a dedicated robustness benchmark tailored explicitly to HOI detection, covering both pixel-level and semantic-level corruptions, is urgently required to put their resilience to the test.
@@
In this paper, we introduce the Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse forms of data corruption. These were designed to simulate real-world conditions. The HICO-DET Chao et al. (2018) and V-COCO Gupta & Malik (2015) datasets served as the starting point. We include robustness-focused metrics, such as the Mean Robustness Index (MRI) Hendrycks & Dietterich (2019) and our novel Composite Robustness Index (CRI). Through our evaluations of existing HOI detection methods on the newly proposed robustness test set, we observe that their performance is less than optimal. This finding underscores the need to develop more robust HOI detection models.
@@
To mitigate the issue of performance degradation, we propose our Semantic-Aware Masking-based Progressive Learning (SAMPL) method. Specifically, we utilize SAM Kirillov et al. (2023) to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying size ratios. This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process.
@@
Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups. Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions. These corruptions simulate realistic semantic and structural disturbances specifically encountered in practical applications.
2 Related Work
@@
Human-Object Interaction Detection. Human-object interaction (HOI) detection Yao & Fei-Fei (2010); Fang et al. (2018); Gupta et al. (2019); Li et al. (2019); Wan et al. (2019); Park et al. (2018); Xu et al. (2020); Yu et al. (2018); Zhang et al. (2015); Zhou et al. (2020) aims to automatically identify and interpret relationships between humans and objects, typically divided into two categories: two-stage and one-stage approaches.
@@
Two-stage HOI detection methods Qi et al. (2018); Xu et al. (2019); Liang et al. (2021); Zhou & Chi (2019) first detect humans and objects independently before classifying interactions. Early CNN-based approaches Xu et al. (2019); Zhou & Chi (2019); Qi et al. (2018) mainly utilized spatial contexts for interaction modeling. Later graph-based methods Zhou & Chi (2019); Xu et al. (2019); Ulutan et al. (2020); Wang et al. (2021); Zhang et al. (2021b) improved relational reasoning explicitly. Subsequent methods further enhanced precision via pose-aware representations Wan et al. (2019), transformer-based unpaired modeling Zhang et al. (2022a), and integration-decomposition strategies Li et al. (2020).
One-stage HOI detection methods Yuan et al. (2023); Zhang et al. (2023) predict human–object interactions in a single forward pass, improving efficiency. Early approaches rely on point-level matching (e.g. PPDM Liao et al. (2020)) or union-level ROI features (e.g. UnionDet Kim et al. (2020)). Later methods refine localization via glance-and-gaze attention (e.g. Zhong et al. (2021)). Transformer-based pipelines further enrich relational context, and methods like RLIPv2 Yuan et al. (2023) combine vision and language modalities for enhanced semantic generalization.
Recent trends explore generative priors and robust-aware training. DIFFUSIONHOI Li et al. (2024) integrates a text-to-image diffusion model to derive relation-aware prompts, while HOI-IDiff Hui et al. (2025) recasts each human–object prediction as an “HOI image” and generates it with a customized diffusion process. UAHOI Chen et al. (2024a) explicitly models prediction uncertainty and incorporates it into the objective, yielding more stable interaction predictions. Although promising, these approaches are still evaluated on clean benchmarks. In contrast, our work introduces RoHOI, a dedicated benchmark of $20$ corruption types, and proposes SAMPL to strengthen robustness under perturbation.
@@
Visual Robustness Benchmarks. Robustness benchmarks have been widely established to assess visual model resilience across various tasks Chattopadhyay et al. (2021); Wang et al. (2024); Sarkar et al. (2024); Schiappa et al. (2024); An et al. (2024); Liu et al. (2024b); Yi et al. (2024); Kamann & Rother (2020); Ma et al. (2024b); Hao et al. (2024); Xie et al. (2025); Hao et al. (2025); Liu et al. (2024c); Michaelis et al. (2019); Zeng et al. (2024), including image classification (ARES-Bench Liu et al. (2024a)), visual question answering Ishmam et al. (2024), common corruption evaluation (ImageNet-C Hendrycks & Dietterich (2019)), attribute editing robustness (ImageNet-E Li et al. (2023)), one-shot skeleton-based action recognition under occlusion Peng et al. (2023), and robust document layout analysis (RoDLA Chen et al. (2024c)). However, robustness benchmarks explicitly designed for HOI detection have not been explored. In this work, we address the uncharted robust HOI detection area by constructing the new RoHOI benchmark and by putting forward the SAMPL method to achieve more reliable HOI detection performance.
3 Benchmark
@@
Corruptions. To rigorously assess the robustness of HOI detection, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under real-world corruptions. Existing benchmarks such as ImageNet-C Hendrycks & Dietterich (2019) primarily cover low-level pixel disturbances like brightness or blur. In contrast, RoHOI explicitly includes semantic-level corruptions specifically tailored to challenge spatial and relational reasoning critical for HOI detection. RoHOI explicitly introduces 20 corruptions derived from HICO-DET and V-COCO images, categorized into four groups: optical system artifacts such as motion blur and glass blur; sensor, compression, and transmission distortions such as packet loss and JPEG artifacts; environmental effects including exposure imbalance and occlusion; and geometric or structural disruptions such as screen cracks, elastic deformation, and perspective distortion. The corruptions range from pixel-level noise to semantic-level challenges that affect spatial alignment and interaction reasoning. Each corruption is implemented with five severity levels to support detailed analysis under progressively more difficult conditions.
@@
Datasets. To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO. HICO-DET Wang et al. (2020) contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories. Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios. V-COCO Gupta & Malik (2015) is a subset of MS-COCO Lin et al. (2014) with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes. Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions. Evaluating these datasets under corruption provides a systematic robustness assessment beyond clean-image performance.
@@
Baselines. To systematically assess robustness, we benchmark representative two-stage and one-stage HOI detectors spanning transformer pipelines and vision–language training paradigms. For the two-stage family, we include UPT Zhang et al. (2022a), which detects humans and objects before interaction classification using a unary–pairwise transformer, and GEN-VLKT Liao et al. (2022), which adopts a guided-embedding association design and a visual–linguistic knowledge transfer (VLKT) strategy that transfers CLIP text embeddings via initialization and a mimic loss rather than performing generic VLM pretraining. For the one-stage family, we evaluate QPIC Tamura et al. (2021) as a query-based pairwise transformer detector, CDN Zhang et al. (2021a) as an end-to-end cascade disentangling network, and DiffHOI Yang et al. (2023) which injects text-to-image diffusion priors to enrich relational cues. We further include FGAHOI Ma et al. (2024a) with fine-grained anchors, QAHOI Chen & Yanai (2023) with query-based anchors, MUREN Kim et al. (2023) with multiplex relational context learning, and SOV-STG Chen et al. (2023) with subject–object–verb decoding guided by specific-target denoising. Finally, we consider RLIPv2 Yuan et al. (2023), which incorporates region-language alignment for fine-grained interaction reasoning. Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning. Our benchmark systematically evaluates these models across diverse corruption types and severity levels, providing a comprehensive assessment of HOI detection robustness.
@@
Evaluation Metrics. We evaluate the clean test set of HICO-DET Wang et al. (2020) using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare (<<10 instances), and non-rare ($\geq$10 instances) categories Yuan et al. (2023). For the clean test set of V-COCO Gupta & Malik (2015), we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$. To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI). MRI, following the definition in prior work Hendrycks & Dietterich (2019), measures average performance under corruptions but overlooks stability across severity levels. To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
@@
$\mathrm{MRI}\;=\;\frac{1}{C}\sum_{c=1}^{C}\left(\frac{1}{L_{c}}\sum_{l=1}^{L_{c}}M_{c,l}\right).$ (1)
While MRI provides an overall robustness estimate, it does not capture variability across severities.
Composite Robustness Index (CRI) counterbalances this by penalizing instability and normalizing to clean-set performance:
$\mathrm{CRI}\;=\;\frac{1}{C}\sum_{c=1}^{C}\left(\frac{\overline{M}_{c}}{M_{\text{clean}}}\cdot\frac{1}{\log\!\big(1+\sigma_{c}\big)+1}\right).$ (2)
Conventions for MRI/CRI. $C$ is the number of corruption types; $L_{c}$ is the number of severity levels for corruption $c$. $M_{c,l}$ denotes the task metric evaluated at level $l$ of corruption $c$; $\overline{M}_{c}$ and $\sigma_{c}$ are, respectively, the mean and standard deviation of $\{M_{c,l}\}_{l=1}^{L_{c}}$; $M_{\text{clean}}$ is the clean-set score. Unless otherwise specified, we instantiate $M$ as mAP (Full) on HICO-DET and $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO.
By combining absolute accuracy with stability across severities, CRI favors models that retain consistent performance under progressively challenging corruptions, complementing the mean-focused view provided by MRI.
4 Method
In this section, we present our novel Semantic-Aware Masking-based Progressive Learning (SAMPL), aiming to enhance model robustness by progressively guiding models from holistic feature reliance towards robust contextual inference. Unlike standard augmentation methods, SAMPL employs adaptive semantic masking guided by SAM Kirillov et al. (2023), dynamically adjusting masking severity based on validation performance. This structured semantic-level masking strategy explicitly promotes robust contextual reasoning by training models to infer interactions from progressively restricted semantic regions, thus effectively improving general robustness against a variety of realistic disturbances. Moreover, SAMPL’s structured progressive curriculum inherently regularizes feature representation, effectively reducing overfitting to dataset biases, thereby further improving generalization under unforeseen corruptions.